# Nyström Kernel Stein Discrepancy

### v2

@@ Appendix B Auxiliary results
$r\left(C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\right)=\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\right)}{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p}}C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}}\leq\frac{\frac{\operatorname{tr}\left(C_{\mathbb{Q},{h_{p}}}\right)}{\lambda}}{\frac{1}{2}},$ (B.41)
we can take $m\geq\max\left\{\frac{2}{\lambda}\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right),\log(1/\delta)\right\}$. To sum up, we got by (B.39) and (B.40) that, conditioned on the Nyström selection, with probability of at least $1-\delta$ it holds that
@@
$\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},h_{p}}-\tilde{C}_{\mathbb{Q},h_{p},m}\right)C_{\mathbb{Q},h_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}^{2}\lesssim\max\left\{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{\lambda m},\frac{\log(1/\delta)}{m}\right\}.$ (B.42)
@@
We remove the conditioning by taking the expectation over all Nyström selections. For $m\gtrsim\max\left\{\frac{\operatorname{tr}\left(C_{\mathbb{Q},h_{p}}\right)}{\lambda},\log(1/\delta)\right\}$, we have that $\beta(\lambda)<1$, which, together with (B.29), implies the stated result. ∎
@@
where in (a) we use that $\frac{x^{n}}{n!}<e^{x}-1$ holds for all $n,x>0$, (b) follows from the definition of the sub-exponential Orlicz norm. ∎
The next lemma shows that $h_{p}\left(\cdot,{X}\right)$ and the “whitened” random variable $C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)$ enjoy sub-Gaussian properties in terms of their respective $\mathcal{H}_{h_{p}}$ norms.
@@
Let $\mathcal{H}$ be a separable Hilbert space, $Y\sim\mathbb{Q}\in\mathcal{M}_{1}^{+}(\mathcal{H})$, and $A\in\mathcal{L}(\mathcal{H})$ invertible, self-adjoint, and positive. Assume that $Y$ is sub-Gaussian, in other words $\left\|{\left\langle{Y,u}\right\rangle_{\mathcal{H}}}\right\|_{\psi_{2}}\lesssim\left\|{\left\langle{Y,u}\right\rangle_{\mathcal{H}}}\right\|_{L_{2}(\mathbb{Q})}$ holds for all $u\in\mathcal{H}$. Then
@@
Specifically, with Assumption 1, choosing $A:=I$ and $Y:=h_{p}\left(\cdot,{X}\right)$, and $A:=C^{-1}_{\mathbb{Q},h_{p},\lambda}$ ($\lambda>0$) and $Y:=h_{p}\left(\cdot,{X}\right)$, respectively, it holds that
$\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}<\infty,$ and $\left\|{\left\|{C_{\mathbb{Q},h_{p},\lambda}^{-1/2}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}^{2}\lesssim\operatorname{tr}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1}C_{\mathbb{Q},h_{p}}\right)<\infty,$ (B.46)
that is, both l.h.s. are sub-Gaussian.
@@
The details are as follows. (a) uses Lemma D.2(4), Parseval’s identity yields (b), and the triangle inequality implies (c). (d) holds by Lemma D.2(4). For (e), let $u_{i}=A^{1/2}e_{i}$ and observe that
@@
With $A:=I$ and $Y:=h_{p}\left(\cdot,{X}\right)$, we obtain $\left\|{\left\|{h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\lesssim\operatorname{tr}\left(\mathbb{E}\left(h_{p}\left(\cdot,{X}\right)\otimes h_{p}\left(\cdot,{X}\right)\right)\right)=\operatorname{tr}(C_{\mathbb{Q},h_{p}})<\infty$, which is the second statement. The last part follows from considering $A:=C^{-1}_{\mathbb{Q},h_{p},\lambda}$ and $Y:=h_{p}\left(\cdot,{X}\right)$; the invertibility of $C_{\mathbb{Q},h_{p},\lambda}$ guarantees the well-definedness of the $u_{i}$-s ($i\in I$).
@@
Let $c>0$ be an absolute constant. As $X_{1}$ is sub-Gaussian, by Vershynin [2018, Proposition 2.5.2], there exists $K_{1}\leq c\left\|{X_{1}}\right\|_{\psi_{2}}$ such that $\mathbb{P}(|X_{1}|\geq t)\leq 2e^{-\frac{t^{2}}{K_{1}^{2}}}$ for all $t\geq 0$. Let $u=\sqrt{K_{1}^{2}(\log(2n)+t)}$
@@
Appendix C Derivation of (9)
We detail the derivation of (9).$(iii)$ in the following. Indeed, for $\mathbf{f}\in\mathcal{H}_{k}^{d}$ and $\bm{\alpha}\in\mathbb{R}^{d}$, we obtain

### v4

@@ Appendix B AUXILIARY RESULTS
$\stackrel{{\scriptstyle(c)}}{{\leq}}\lambda\left\|{C_{\tilde{\mathbb{Q}}_{m},\tilde{h}_{p},\lambda}^{-1/2}C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}$ (B.43)
where (a) follows from Rudi et al. [2015, Proposition 3] with $X:=C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}$ therein. (b) is by Sterge and Sriperumbudur [2022, Lemma A.7(iv)]. Lemma B.6(5) with $C:=C_{\tilde{\mathbb{Q}}_{m},h_{p},\lambda}^{-1/2}$, $D:=C_{\tilde{\mathbb{Q}}_{m},\tilde{h}_{p},\lambda}^{-1/2}$, and $X:=C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}$ yields (c), as we obtain $C^{*}C=C_{\tilde{\mathbb{Q}}_{m},h_{p},\lambda}^{-1}\preccurlyeq C_{\tilde{\mathbb{Q}}_{m},\tilde{h}_{p},\lambda}^{-1}=D^{*}D$; the positive definite relationship holding by the following chain of inequalities
[…]
$\stackrel{{\scriptstyle(e)}}{{\iff}}0\preccurlyeq\mu_{h_{p}}\left(\tilde{\mathbb{Q}}_{m}\right)\otimes\mu_{h_{p}}\left(\tilde{\mathbb{Q}}_{m}\right),$ (B.45)
which is true as the r.h.s. is a positive operator. In (d), we subtract $\lambda I$ from both sides. (e) follows from subtracting $C_{\tilde{\mathbb{Q}}_{m},h_{p}}$ and by multiplying with $-1$.
Applying the second inequality in the statement of Rudi et al. [2015, Proposition 7] to (B.42) (we specialize $A:=C_{\tilde{\mathbb{Q}}_{m},\tilde{h}_{p}}$ and $B:=C_{\mathbb{Q},\bar{h}_{p}}$ therein), we obtain
@@
$\lambda\left\|{C_{\tilde{\mathbb{Q}}_{m},\tilde{h}_{p},\lambda}^{-1/2}C_{\mathbb{Q},\bar{h}_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\leq\frac{\lambda}{1-\beta(\lambda)},$ (B.46)
when $\beta(\lambda)<1$. The combination of (B.42) and (B.46) yields the first stated claim.
Step 2.
It remains to show that $\beta(\lambda)<1$ holds with high probability. Let us introduce the shorthands $\tilde{\mu}_{h_{p}}=\mu_{h_{p}}\left(\tilde{\mathbb{Q}}_{m}\right)=\frac{1}{m}\sum_{i\in[m]}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)$ and $\mu_{h_{p}}=\mu_{h_{p}}(\mathbb{Q})$. Notice that we have
$C_{\tilde{\mathbb{Q}}_{m},\tilde{h}_{p}}=C_{\tilde{\mathbb{Q}}_{m},\bar{h}_{p}}-\left[\tilde{\mu}_{h_{p}}-\mu_{h_{p}}\right]\otimes\left[\tilde{\mu}_{h_{p}}-\mu_{h_{p}}\right],$ (B.47)
which is verified by using the linearity of tensor products and by using that
$C_{\tilde{\mathbb{Q}}_{m},\bar{h}_{p}}=\frac{1}{m}\sum_{i\in[m]}\bar{h}_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\otimes\bar{h}_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right).$ (B.48)
Instead of showing that $\beta(\lambda)<1$, we will show that the following stronger requirement can be satisfied:
$\beta(\lambda)\stackrel{{\scriptstyle(a)}}{{\leq}}\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},\bar{h}_{p}}-C_{\tilde{\mathbb{Q}}_{m},\tilde{h}_{p}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}$ (B.49)
=$\stackrel{{\scriptstyle(b)}}{{=}}\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},\bar{h}_{p}}-C_{\tilde{\mathbb{Q}}_{m},\bar{h}_{p}}+\left[\tilde{\mu}_{h_{p}}-\mu_{h_{p}}\right]\otimes\left[\tilde{\mu}_{h_{p}}-\mu_{h_{p}}\right]\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}$ (B.50)
$\stackrel{{\scriptstyle(c)}}{{\leq}}\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},\bar{h}_{p}}-C_{\tilde{\mathbb{Q}}_{m},\bar{h}_{p}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}+$ (B.51)
$\quad+\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(\left[\tilde{\mu}_{h_{p}}-\mu_{h_{p}}\right]\otimes\left[\tilde{\mu}_{h_{p}}-\mu_{h_{p}}\right]\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}$ (B.52)
=$\stackrel{{\scriptstyle(d)}}{{=}}\underbrace{\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},\bar{h}_{p}}-C_{\tilde{\mathbb{Q}}_{m},\bar{h}_{p}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}}_{=:T_{1}}+\underbrace{\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(\tilde{\mu}_{h_{p}}-\mu_{h_{p}}\right)}\right\|_{\mathcal{H}_{h_{p}}}^{2}}_{=:T_{2}}<1.$ (B.53)
In (a), we use that the spectral radius is bounded by the operator norm. (b) uses (B.47) and (c) holds by the triangle inequality. Lemma B.4 and Lemma C.1 applied to the second term yield (d).
  • First term ($T_{1}$). We will bring ourselves into the setting of Koltchinskii and Lounici [2017, Theorem 9] (recalled in Theorem C.3). First, we condition on the Nyström selection and define the centered random variables $\eta_{i_{j}}=C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{j}}\right)-\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)\right)$ ($=C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\bar{h}_{p}(\cdot,\tilde{\mathbf{x}}_{j})$) ($j\in[m]$), which satisfy the sub-Gaussian assumption. Indeed, let $u\in\mathcal{H}_{h_{p}}$ be arbitrary, and $v=C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}u\in\mathcal{H}_{h_{p}}$; the invertibility of $C_{\mathbb{Q},\bar{h}_{p},\lambda}$ guarantees the well-definedness of $v$. With this notation, for any $j\in[m]$,
    $\left\|{\left\langle{\eta_{i_{j}},u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(a)}}{{=}}\left\|{\left\langle{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\bar{h}_{p}(\cdot,\tilde{\mathbf{x}}_{j}),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(b)}}{{=}}\left\|{\left\langle{\bar{h}_{p}(\cdot,\tilde{\mathbf{x}}_{j}),C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}$ (B.54)
    =$\stackrel{{\scriptstyle(c)}}{{=}}\left\|{\left\langle{\bar{h}_{p}(\cdot,\tilde{\mathbf{x}}_{j}),v}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\stackrel{{\scriptstyle(d)}}{{\lesssim}}\underbrace{\left\|{\left\langle{\bar{h}_{p}(\cdot,\tilde{\mathbf{x}}_{j}),v}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}}_{(\dagger)}\stackrel{{\scriptstyle(e)}}{{=}}\left\|{\left\langle{\bar{h}_{p}(\cdot,\tilde{\mathbf{x}}_{j}),C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}$ (B.55)
    =$\stackrel{{\scriptstyle(f)}}{{=}}\left\|{\left\langle{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\bar{h}_{p}(\cdot,\tilde{\mathbf{x}}_{i_{j}}),u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}\stackrel{{\scriptstyle(g)}}{{=}}\left\|{\left\langle{\eta_{i_{j}},u}\right\rangle_{\mathcal{H}_{h_{p}}}}\right\|_{L_{2}(\mathbb{Q})}<\infty.$ (B.56)
    (a) is the definition of the $\eta_{i_{j}}$-s, (b) uses the self-adjointness of $C_{\mathbb{Q},\bar{h}_{p},\lambda}$, and (c) follows from the definition of $v$. The sub-Gaussian assumption implies (d), (e) again follows from the definition of $v$, and (f) is implied by the self-adjointness of $C_{\mathbb{Q},\bar{h}_{p},\lambda}$. Inserting the definition of $\eta_{i_{j}}$ in (g) proves their sub-Gaussianity by using that ()<($\dagger$)$\infty$( † ) < ∞ according to Assumption 1 and as the derivation afterwards only involved equalities.
    Let $A=C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}C_{\mathbb{Q},\bar{h}_{p}}C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}$. Theorem C.3 yields that, conditioned on the Nyström selection, it holds with probability at least $21-\delta/$ 2 that
    $\Big{\|}\underbrace{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},\bar{h}_{p}}-C_{\tilde{\mathbb{Q}}_{m},\bar{h}_{p}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}}_{=\frac{1}{m}\sum_{j=1}^{m}\eta_{i_{j}}\otimes\eta_{i_{j}}-\mathbb{E}\left[\eta_{i_{j}}\otimes\eta_{i_{j}}\right]}\Big{\|}_{\text{op}}\lesssim\left\|{A}\right\|_{\mathrm{op}}\max\left(\sqrt{\frac{r\left(A\right)}{m}},\sqrt{\frac{\log(2/\delta)}{m}}\right),$ (B.57)
    provided that $m\geq\max\left\{r\left(A\right),\log(2/\delta)\right\}$, with $r\left(A\right)=\frac{\operatorname{tr}\left(A\right)}{\left\|{A}\right\|_{\mathrm{op}}}$. Using Lemma B.6(2), $A\preccurlyeq I$, hence $\left\|{A}\right\|_{\mathrm{op}}\leq 1$. Moreover by Lemma B.7(3), $r(A)\leq\frac{2\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{\lambda}$, which implies that, with the same probability,
    $\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},\bar{h}_{p}}-C_{\tilde{\mathbb{Q}}_{m},\bar{h}_{p}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}\lesssim\max\left(\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{\lambda m}},\sqrt{\frac{\log(2/\delta)}{m}}\right),$ (B.58)
    holds when $m\geq\max\left\{\frac{2\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{\lambda},\log(2/\delta)\right\}$. Therefore, one can take
    $m\gtrsim\max\left\{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{\lambda},\log(2/\delta)\right\}$ (B.59)
    to get $\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},\bar{h}_{p}}-C_{\tilde{\mathbb{Q}}_{m},\bar{h}_{p}}\right)C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}}\right\|_{\mathrm{op}}<\frac{1}{2}$ holding with probability at least $21-\delta/$ 2.
  • Second term ($T_{2}$). We condition again on the Nyström selection, let $\eta_{i_{j}}=C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\bar{h}_{p}\left(\cdot,\mathbf{x}_{i_{j}}\right)$ for $j\in[m]$, and observe that $\frac{1}{m}\sum_{j\in[m]}\eta_{i_{j}}=C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\left(\tilde{\mu}_{h_{p}}-\mu_{h_{p}}\right)$. The $\eta_{i_{j}}$-s are centered, and, by Lemma B.3, it holds for any $j\in[m]$ that
    $\left\|{\left\|{\eta_{i_{j}}}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}^{2}\lesssim\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1}C_{\mathbb{Q},\bar{h}_{p}}\right),$ (B.60)
    that is, the $\left\|{\eta_{i_{j}}}\right\|_{\mathcal{H}_{h_{p}}}$-s are sub-Gaussian. Hence, by Lemma C.2(3), they are sub-exponential, and, by Lemma B.2, they satisfy the Bernstein condition with $\sigma,B\lesssim\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1}C_{\mathbb{Q},\bar{h}_{p}}\right)}$. Therefore, application of Theorem C.4 yields that, conditioned on the Nyström choice, it holds with probability at least $21-\delta/$ 2 that
    $\left\|{\frac{1}{m}\sum_{j=1}^{m}\eta_{i_{j}}}\right\|_{\mathcal{H}_{h_{p}}}$ $\lesssim\frac{\sqrt{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1}C_{\mathbb{Q},\bar{h}_{p}}\right)}\log(4/\delta)}{m}+\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1}C_{\mathbb{Q},\bar{h}_{p}}\right)\log(4/\delta)}{m}}$ (B.61)
    $\stackrel{{\scriptstyle(a)}}{{\lesssim}}\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1}C_{\mathbb{Q},\bar{h}_{p}}\right)\log(4/\delta)}{m}}\stackrel{{\scriptstyle(b)}}{{\leq}}\sqrt{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)\log(4/\delta)}{\lambda m}}$ (B.62)
    where in (a), we assume that $m\geq\log(4/\delta)$ and notice that this condition implies that the first term is smaller than the second term. Lemma B.7(1) yields (b). The obtained bound means that choosing $m\gtrsim\max\left\{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{\lambda},1\right\}\log(4/\delta)$ guarantees that $\left\|{\frac{1}{m}\sum_{j=1}^{m}\eta_{i_{j}}}\right\|_{\mathcal{H}_{h_{p}}}^{2}<\frac{1}{2}$ holds with probability at least $21-\delta/$ 2.
@@
As a final step, we observe that $\log(2/\delta)<\log(4/\delta)$ and $\log(4/\delta)>1$, which, by union bound, shows that, for $m\gtrsim\max\left\{\frac{\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p}}\right)}{\lambda},1\right\}\log(4/\delta)$, it holds with probability at least $1-\delta$ that $\beta(\lambda)<1$. We lift the conditioning by integrating over all Nyström selections. ∎
@@
where in (a) we use that $\frac{x^{n}}{n!}<e^{x}-1$ holds for all $n,x>0$, and (b) follows from the definition of the sub-exponential Orlicz norm. ∎
The next lemma shows that $\bar{h}_{p}\left(\cdot,{X}\right)$ and the “whitened” random variable $C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\bar{h}_{p}\left(\cdot,{X}\right)$ enjoy sub-Gaussian properties in terms of their respective $\mathcal{H}_{h_{p}}$ norms.
@@
Let $\mathcal{H}$ be a separable Hilbert space, $Y\sim\mathbb{Q}\in\mathcal{M}_{1}^{+}(\mathcal{H})$, and $A\in\mathcal{L}(\mathcal{H})$ invertible, and positive. Assume that $Y$ is sub-Gaussian, in other words $\left\|{\left\langle{Y,u}\right\rangle_{\mathcal{H}}}\right\|_{\psi_{2}}\lesssim\left\|{\left\langle{Y,u}\right\rangle_{\mathcal{H}}}\right\|_{L_{2}(\mathbb{Q})}$ holds for all $u\in\mathcal{H}$. Then
@@
Specifically, with Assumption 1, choosing $A:=I$ and $Y:=\bar{h}_{p}\left(\cdot,{X}\right)$, and $A:=C^{-1}_{\mathbb{Q},\bar{h}_{p},\lambda}$ ($\lambda>0$) and $Y:=\bar{h}_{p}\left(\cdot,{X}\right)$, respectively, it holds that
$\left\|{\left\|{\bar{h}_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}<\infty,$ and $\left\|{\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\bar{h}_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}^{2}\lesssim\operatorname{tr}\left(C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1}C_{\mathbb{Q},\bar{h}_{p}}\right)<\infty,$ (B.66)
that is, both $\left\|{\bar{h}_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}$ and $\left\|{C_{\mathbb{Q},\bar{h}_{p},\lambda}^{-1/2}\bar{h}_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}$ are sub-Gaussian.
@@
The details are as follows. (a) uses Lemma C.2(4), Parseval’s identity yields (b), and the triangle inequality implies (c). (d) holds by Lemma C.2(4). For (e), let $u_{i}=A^{1/2}e_{i}$ and observe that
@@
With $A:=I$ and $Y:=\bar{h}_{p}\left(\cdot,{X}\right)$, we have $\left\|{\left\|{\bar{h}_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}}\right\|_{\psi_{2}}\lesssim\operatorname{tr}\left(\mathbb{E}\left(\bar{h}_{p}\left(\cdot,{X}\right)\otimes\bar{h}_{p}\left(\cdot,{X}\right)\right)\right)=\operatorname{tr}(C_{\mathbb{Q},\bar{h}_{p}})<\infty$, which is the second statement. The last part follows from considering $A:=C^{-1}_{\mathbb{Q},\bar{h}_{p},\lambda}$ and $Y:=\bar{h}_{p}\left(\cdot,{X}\right)$; the invertibility of $C_{\mathbb{Q},\bar{h}_{p},\lambda}$ guarantees the well-definedness of the $u_{i}$-s ($i\in I$). ∎
@@
Let $c>0$ be an absolute constant. As $X_{1}$ is sub-Gaussian, by Vershynin [2018, Proposition 2.5.2], there exists $K_{1}\leq c\left\|{X_{1}}\right\|_{\psi_{2}}$ such that $\mathbb{P}(|X_{1}|\geq t)\leq 2e^{-\frac{t^{2}}{K_{1}^{2}}}$ for all $t\geq 0$. Let $u=\sqrt{K_{1}^{2}(\log(2n)+t)}$. Then
@@
The following result shows that positive operators share some well-known properties of positive (semi-)definite matrices; we refer to Bhatia [2007] for the related matrix cases.
Lemma B.6 (Properties of positive operators).
Let $\mathcal{H}$ be a Hilbert space and assume $A,B\in\mathcal{L}\left(\mathcal{H}\right)$ are positive and invertible. Then, the following hold.
  1. 1.
    If $A\preccurlyeq B$, then $X^{*}AX\preccurlyeq X^{*}BX$ for any $X\in\mathcal{L}\left(\mathcal{H}\right)$.
  2. 2.
    If $A\preccurlyeq B$, then $B^{-1/2}AB^{-1/2}\preccurlyeq I$.
  3. 3.
    If $B\preccurlyeq I$, then $B^{-1}\succcurlyeq I$.
  4. 4.
    If $A\preccurlyeq B$, then $A^{-1}\succcurlyeq B^{-1}$.
  5. 5.
    If $C^{*}C\preccurlyeq D^{*}D$, then $\left\|{CX}\right\|_{\mathrm{op}}\leq\left\|{DX}\right\|_{\mathrm{op}}$ for any $C,D,X\in\mathcal{L}\left(\mathcal{H}\right)$.
Proof.
  1. 1.